OpenStudy (jjuden):

Which number is halfway between 1/3 and 1/2?

OpenStudy (jjuden):

a 2/12 b 3/12 c 5/12 d 7/12

OpenStudy (isaiah.feynman):

I think you can figure this one out. After all you answered a question on the pythagorean theorem

OpenStudy (jjuden):

i have my week spots in math

OpenStudy (jjuden):

and this is one of them

OpenStudy (isaiah.feynman):

Numbers?

OpenStudy (jjuden):

i got it nm guys

OpenStudy (ybarrap):

$$ 1/3 = 4/12\\ 1/2=6/12 $$ So now, compare numerators.

OpenStudy (jjuden):

5/12 is what got

OpenStudy (isaiah.feynman):